OpenStudy (anonymous):

If a is rational and b is irrational, is ab necessarily irrational? (Careful!)

myininaya (myininaya):

we need a contradiction

OpenStudy (anonymous):

may be, not really

OpenStudy (anonymous):

u can use that in one case

myininaya (myininaya):

b=baa^(-1) is irrational b=aba^(-1)=(ab)(a^(-1)) a is rational => a^(-1) is rational (assuming a doesn't =0) but for b to be irrational that means ab has to be irrational

OpenStudy (nikita2):

if a = 0 then ab is rational. Assume a not= 0 and ab is rational then ab/a is rational too, so b = ab/a has to be rational.

myininaya (myininaya):

but 0 is rational 0*b is rational

OpenStudy (anonymous):

hey nikita do u have this book? ny Michael Spivak?

OpenStudy (anonymous):

by*

OpenStudy (nikita2):

definitely no )

OpenStudy (anonymous):

okay ur approach is right

OpenStudy (anonymous):

two cases: if a = 0 if a is not 0

OpenStudy (nikita2):

wat is title of this book?

OpenStudy (anonymous):

Calculus by Michael Spivak

OpenStudy (anonymous):

thats right, its a good book

OpenStudy (nikita2):

Are u really want to learn it ?

OpenStudy (anonymous):

what do u mean?

OpenStudy (nikita2):

I mean calculus and this boot in particular